# On orthogonality graphs of Okubo algebras

### v1

@@ 1 Introduction
$\mathbb{P}(X)=\{[x]=\mathbb{F}x\;|\;x\in X\setminus\{0\}\}.$
@@
The aim of our study is to investigate relation graphs of an important class of non-associative algebras, namely, Okubo algebras. Okubo algebras were first defined in Okubo’s paper [17] and have subsequently been extensively studied in numerous papers by Okubo and Osborn [16, 17, 18], as well as by Elduque [3, 4, 5, 6, 7, 8], who also co-authored with Myung [9, 10] and Pérez [11, 12]. According to [7, Theorem 2.9], Okubo algebras, along with forms of para-Hurwitz algebras, constitute the class of symmetric composition algebras. Furthermore, any finite-dimensional flexible composition algebra is either unital (i.e., it is a Hurwitz algebra) or symmetric, see [11, Theorem 3.2].
In our previous paper [20] commuting graphs of Okubo algebras have been considered. We have particularly focused on the case of the pseudo-octonion algebra over a field $\mathbb{F}$, $\operatorname{char}\mathbb{F}\neq 3$, which contains a primitive cube root of unity, and also on the case of the real division Okubo algebra — the so-called real pseudo-octonions. Besides, by [20, Corollary 4.6], the distance between any two idempotents in the commuting graph of an arbitrary Okubo algebra is at most $2$, that is, the intersection of their centralizers is nonzero.
@@
Section 3 is devoted to orthogonality graphs of Okubo algebras with isotropic norm over arbitrary fields. Proposition 3.1 states that any zero divisor $x$ satisfying $\operatorname{n}(x,x*x)\neq 0$ generates a connected component of diameter $1$ which consists of only two vertices. If an Okubo algebra contains a nonzero element $x$ such that $\operatorname{n}(x,x*x)=0$, then it necessarily has a nonzero idempotent [14, (34.10)]. Theorem 3.15 describes the connected components of the orthogonality graph which are formed by such elements. Their number and diameters depend on the characteristic of the field $\mathbb{F}$, the existence of a primitive cube root of unity $\omega$ in $\mathbb{F}$, and the parameter $\beta$ which determines the multiplication table 1 of an Okubo algebra with isotropic norm and nonzero idempotents. More specifically, if $\operatorname{char}\mathbb{F}\neq 3$ and $\omega\notin\mathbb{F}$, then these elements form several connected components, and each of them is a star graph, cf. Lemma 3.9. Otherwise, all elements such that $\operatorname{n}(x,x*x)=0$ belong to the same connected component whose diameter equals $5$ if the Okubo algebra is split, and $3$ otherwise, see Lemmas 3.10 and 3.14 and Example 3.13.
@@
In Section 5 we establish a connection between the orthogonality graphs of Okubo algebras over fields which contain the primitive cube root of unity $\omega$ and the orthogonality graphs of matrix rings. In particular, it is shown that the connected component that consists of the elements satisfying $\operatorname{n}(x,x*x)=0$ is isomorphic to the subgraph of the orthogonality graph of $3\times 3$ matrices on the set of nilpotent matrices, which was studied in detail by Bakhadly, Guterman, and Markova [1]. This gives an alternative proof of Theorem 3.15 in the case when $\operatorname{char}\mathbb{F}\neq 3$ and $\omega\in\mathbb{F}$.
@@ 2.1 Construction of Okubo algebras
We use the definition of an Okubo algebra given in [4, Section 1]. Assume first that the characteristic of the field $\mathbb{F}$ is distinct from $3$, and $\mathbb{F}$ contains a primitive cube root of unity $\omega$. In particular, the second condition is satisfied if $\mathbb{F}$ is algebraically closed. We set
@@
on the algebra $P_{8}(\mathbb{F})$. As noted in [4, p. 101], this definition makes sense even when $\operatorname{char}\mathbb{F}=2$. Indeed, by [13, p. 1027, Lemma], for any matrix $x\in M_{3}(\mathbb{F})$ the identity
@@
Let $(\mathcal{A},+,*)$ be an arbitrary (possibly non-unital and non-associative) algebra over a field $\mathbb{F}$. Assume that $\mathcal{A}$ is equipped with a strictly nondegenerate quadratic form $\operatorname{n}(\cdot)$, i.e., the associated symmetric bilinear form $\operatorname{n}(a,b)=\operatorname{n}(a+b)-\operatorname{n}(a)-\operatorname{n}(b)$ is nondegenerate on $\mathcal{A}$. Then $\mathcal{A}$ is called a composition algebra if the quadratic form $\operatorname{n}(\cdot)$ admits composition, i.e., $\operatorname{n}(a*b)=\operatorname{n}(a)\operatorname{n}(b)$ for all $a,b\in\mathcal{A}$. Here $\operatorname{n}(\cdot)$ is called the norm.
@@
By [10, Lemma 1.1], the algebra $P_{8}(\mathbb{F})$ is a symmetric composition algebra. Note that Eq. (2.2) and the identity $\operatorname{tr}(xy)=\operatorname{tr}(yx)$ imply that
@@
$\widetilde{P}_{8}(\mathbb{R})=\big\{x\in M_{3}(\mathbb{C})\>|\>x^{*}=x,\>\operatorname{tr}x=0\big\}.$
@@
for all $x\in\mathcal{O}$, see [14, (34.3)]. Together with Theorem 2.3(2), this implies that the subalgebra generated by an element $x$ coincides with the linear span of $x$ and xxx*x.
@@ 2.2 Multiplication table of Okubo algebras with isotropic norm
A special role in the study of Okubo algebras is played by algebras of the form $\mathcal{O}_{\alpha,\beta}$. Given nonzero $\alpha,\beta\in\mathbb{F}$, the multiplication table of $\mathcal{O}_{\alpha,\beta}$ is described by Table 1, while basis elements satisfy the conditions [8, pp. 3-4]:
@@
Theorem 2.7 ([4, Theorem 7], [8, Theorem 4.2]).
@@
Theorem 2.8 ([5, Theorem 3.18], [7, Theorem 5.9]).
@@
Definition 2.10 ([6, Definition 22], [7, Definition 6.2]).
@@
Theorem 2.11 ([2, Proposition 9.9 and Theorem 9.13], [6, Lemma 21], [7, Corollary 6.5]).
@@
The relationship between the quaternionic idempotent and all other idempotents (quadratic and singular) of the split Okubo algebra is described in [7, Proposition 8.7].
It is also known that if $\operatorname{char}\mathbb{F}=3$, then the norm on any Okubo algebra over $\mathbb{F}$ is isotropic [3, Theorem 5.1]. Besides, any Okubo algebra over a finite field is split [8, Theorem 3.2].
The following proposition utilizes the classification of Okubo algebras via central simple algebras of degree $3$ which was obtained in [10]. It should be noted that these results were originally stated for $\operatorname{char}\mathbb{F}\notin\{2,3\}$, however, they still remain valid for $\operatorname{char}\mathbb{F}=2$, see [3, pp. 286–287].
Proposition 2.12 ([10, p. 2486, p. 2501, Propositions 7.3 and 7.4], [12, Corollary 3.6]).
@@ 2.3 Zero divisors in Okubo algebras
  • The orthogonalizer of $a$ is $O_{\mathcal{A}}(a)=\big\{b\in\mathcal{A}\>|\;ab=ba=0\big\}$, i.e., the set of all elements in $\mathcal{A}$ which are orthogonal to $a$.
  • The left annihilator of $a$ is the set $l.\operatorname{Ann}_{\mathcal{A}}(a)=\big\{b\in\mathcal{A}\>|\;ba=0\big\}$.
  • The right annihilator of $a$ is the set $r.\operatorname{Ann}_{\mathcal{A}}(a)=\big\{b\in\mathcal{A}\>|\;ab=0\big\}$.
@@
$x^{\perp}=\{a\;|\;\operatorname{n}(x,a)=0\}.$
Lemma 2.17 ([15, Theorem 3.1, Proposition 3.4]).
@@
$l.\operatorname{Ann}_{\mathcal{O}}(x)$ $=x*\mathcal{O}=\{x*a\;|\;a\in\mathcal{O}\},$
$r.\operatorname{Ann}_{\mathcal{O}}(x)$ $=\mathcal{O}*x=\{a*x\;|\;a\in\mathcal{O}\}.$
Lemma 2.18 ([15, Proposition 3.8]).
@@
Note that [15, Proposition 3.7] also describes the intersection of two left (or two right) annihilators of linearly independent zero divisors in $\mathcal{O}$, but it will not be used in the current paper.
@@ 3 Orthogonality graph of an Okubo algebra
and thus $\operatorname{n}(x,y)=0$. Note that the desired equality can be obtained directly from [15, Proposition 3.7] by using the fact that $z\in(x*\mathcal{O})\cap(y*\mathcal{O})$. ∎
@@
$V^{\prime}_{x}=\big\{[y]=\mathbb{F}y\;|\;y\in O_{\mathcal{O}}(x)\big\}$
@@
$V=\big\{[x]=\mathbb{F}x\;|\;x\in Z(\mathcal{O}),\>\operatorname{n}(x,x*x)=0\big\}$
@@
$V=\big\{[x]=\mathbb{F}x\;|\;x\in Z(\mathcal{O}),\>\operatorname{n}(x,x*x)=0\big\}$
@@ 4 Number of shortest paths
$C_{(\mathcal{S},*,\operatorname{n})}(e)=\big\{x\in\mathcal{S}\>|\;e*x=x*e\big\},$
@@
$\operatorname{Fix}(\tau)=\big\{x\in\mathcal{S}\>|\;\tau(x)=x\big\}.$
Proposition 4.9 ([2, Proposition 9.9 and Theorem 9.13]).
@@
In [7, Theorem 6.3] all possible cases of Petersson algebras associated to the Zorn vector-matrix algebra over a field of characteristic $3$ are described in detail. We will only cite the first item of this statement.
Theorem 4.10 ([7, Theorem 6.3(i)]).
@@
Proposition 4.11 ([6, Proposition 24], [7, Theorem 8.5]).
@@
    $\big\{e+x\;|\;0\neq x\in C_{\mathcal{O}}(e)\cap C_{\mathcal{O}}(e)^{\perp}\big\};$
@@
Let $\tau$ be the automorphism defined in Eq. (4.2) which corresponds to the idempotent $e$. Consider a new multiplication on $\mathcal{O}$ given by Eq. (4.3). Then, by Proposition 4.7, the Okubo algebra $(\mathcal{O},*,\operatorname{n})$ is the Petersson algebra associated to the Hurwitz algebra $(\mathcal{O},\cdot,\operatorname{n})$ with the unity $e$ and its automorphism $\tau$. It follows from [7, Theorem 6.3] that $\tau$ satisfies the conditions of Theorem 4.10. Then $(\mathcal{O},\cdot,\operatorname{n})$ is isomorphic to the Zorn vector-matrix algebra over $\mathbb{F}$, and there exists a canonical basis such that $\tau$ acts on basis vectors according to Eqs. (4.4). It remains to show that if xx=0x*x=0 for some $x\in\mathcal{O}$, then $x\in C_{\mathcal{O}}(e)=\operatorname{span}\{e_{1},e_{2},u_{1},u_{2},v_{1},v_{3}\}$.
@@
Consider $x=(z_{0,2}+z_{1,2}+z_{2,2})-(z_{0,1}+z_{2,1}+z_{1,1})$. Since $x\perp C_{\mathcal{O}}(e)$, the element $e+x$ is a singular idempotent (see also an example after Definition 22 in [6] and [5, Theorem 3.12]). It is shown in Lemma 3.14 that yy=0y*y=0 for all $y\in O_{\mathcal{O}}(x)$. Clearly, this property of the orthogonalizer is preserved under automorphism, so it holds for all elements of type (1).

### v2

@@ 1 Introduction
$\mathbb{P}(X)=\{[x]=\mathbb{F}x\mid x\in X\setminus\{0\}\}.$
@@
Among the directions where relation graphs find particularly important applications, we mention two intertwining questions: the problem of classification of relation preserving mappings and the isomorphism problem, that is, exploring the connection between the isomorphism of algebraic structures and the isomorphism of corresponding relation graphs. For instance, in [1] both these questions were answered for Birkhoff–James orthogonality graphs of smooth normed spaces, and in [5] the isomorphism problem of finite-dimensional formally real simple Jordan algebras was completely solved in terms of graphs induced by Jordan orthogonality. Note also that linear preserver problems are closely connected to Jordan homomorphisms, see a recent survey by Brešar and Zelmanov [3].
The aim of our study is to investigate relation graphs of an important class of non-associative algebras, namely, Okubo algebras. Okubo algebras were first defined in Okubo’s paper [21] and have subsequently been extensively studied in numerous papers by Okubo and Osborn [20, 21, 19], as well as by Elduque [10, 11, 12, 13, 14, 15], who also co-authored with Myung [6, 7] and Pérez [8, 9]. According to [14, Theorem 2.9], Okubo algebras, along with forms of para-Hurwitz algebras, constitute the class of symmetric composition algebras. Furthermore, any finite-dimensional flexible composition algebra is either unital (i.e., it is a Hurwitz algebra) or symmetric, see [8, Theorem 3.2].
In our previous paper [23] commuting graphs of Okubo algebras have been considered. We have particularly focused on the case of the pseudo-octonion algebra over a field $\mathbb{F}$, $\operatorname{char}\mathbb{F}\neq 3$, which contains a primitive cube root of unity, and also on the case of the real division Okubo algebra — the so-called real pseudo-octonions. Besides, by [23, Corollary 4.6], the distance between any two idempotents in the commuting graph of an arbitrary Okubo algebra is at most $2$, that is, the intersection of their centralizers is nonzero.
@@
We remark that our approach to the study of Okubo algebras leads to deeper understanding of their idempotents and the role played by zero divisors satisfying the condition $\operatorname{n}(x,x*x)=0$. In particular, in the case of the split Okubo algebra over a field of characteristic $3$, Corollary 4.15 extends the characterization of singular and quadratic idempotents obtained in [14, Proposition 8.7] to terms of the orthogonalizers of their corresponding zero divisors, and Corollary 4.17 reveals how these zero divisors interact in the orthogonality graph.
@@
Section 3 is devoted to orthogonality graphs of Okubo algebras with isotropic norm over arbitrary fields. Proposition 3.1 states that any zero divisor $x$ satisfying $\operatorname{n}(x,x*x)\neq 0$ generates a connected component of diameter $1$ which consists of only two vertices. If an Okubo algebra contains a nonzero element $x$ such that $\operatorname{n}(x,x*x)=0$, then it necessarily has a nonzero idempotent [17, (34.10)]. Theorem 3.15 describes the connected components of the orthogonality graph which are formed by such elements. Their number and diameters depend on the characteristic of the field $\mathbb{F}$, the existence of a primitive cube root of unity $\omega$ in $\mathbb{F}$, and the parameter $\beta$ which determines the multiplication table 1 of an Okubo algebra with isotropic norm and nonzero idempotents. More specifically, if $\operatorname{char}\mathbb{F}\neq 3$ and $\omega\notin\mathbb{F}$, then these elements form several connected components, and each of them is a star graph, cf. Lemma 3.9. Otherwise, all elements such that $\operatorname{n}(x,x*x)=0$ belong to the same connected component whose diameter equals $5$ if the Okubo algebra is split, and $3$ otherwise, see Lemmas 3.10 and 3.14 and Example 3.13.
@@
In Section 5 we establish a connection between the orthogonality graphs of Okubo algebras over fields which contain the primitive cube root of unity $\omega$ and the orthogonality graphs of matrix rings. In particular, it is shown that the connected component that consists of the elements satisfying $\operatorname{n}(x,x*x)=0$ is isomorphic to the subgraph of the orthogonality graph of $3\times 3$ matrices on the set of nilpotent matrices, which was studied in detail by Bakhadly, Guterman, and Markova [2]. This gives an alternative proof of Theorem 3.15 in the case when $\operatorname{char}\mathbb{F}\neq 3$ and $\omega\in\mathbb{F}$. On the other hand, the results of Section 4 show that the orthogonality graph of nilpotent $3\times 3$ matrices over such a field $\mathbb{F}$ is bigeodetic, see Remark 5.8.
@@ 2.1 Construction of Okubo algebras
We use the definition of an Okubo algebra given in [11, Section 1]. Assume first that the characteristic of the field $\mathbb{F}$ is distinct from $3$, and $\mathbb{F}$ contains a primitive cube root of unity $\omega$. In particular, the second condition is satisfied if $\mathbb{F}$ is algebraically closed. We set
@@
on the algebra $P_{8}(\mathbb{F})$. As noted in [11, p. 101], this definition makes sense even when $\operatorname{char}\mathbb{F}=2$. Indeed, by [16, p. 1027, Lemma], for any matrix $x\in M_{3}(\mathbb{F})$ the identity
@@
Let $(\mathcal{A},+,*)$ be an arbitrary (possibly non-unital and non-associative) algebra over a field $\mathbb{F}$. Assume that $\mathcal{A}$ is equipped with a strictly nondegenerate quadratic form $\operatorname{n}(\cdot)$, i.e., the associated symmetric bilinear form
$\operatorname{n}(a,b)=\operatorname{n}(a+b)-\operatorname{n}(a)-\operatorname{n}(b)$
is nondegenerate on $\mathcal{A}$. Then $\mathcal{A}$ is called a composition algebra if the quadratic form $\operatorname{n}(\cdot)$ admits composition, i.e., $\operatorname{n}(a*b)=\operatorname{n}(a)\operatorname{n}(b)$ for all $a,b\in\mathcal{A}$. Here $\operatorname{n}(\cdot)$ is called the norm.
@@
By [7, Lemma 1.1], the algebra $P_{8}(\mathbb{F})$ is a symmetric composition algebra. Note that Eq. (2.2) and the identity $\operatorname{tr}(xy)=\operatorname{tr}(yx)$ imply that
@@
$\widetilde{P}_{8}(\mathbb{R})=\big\{x\in M_{3}(\mathbb{C})\mid x^{*}=x,\>\operatorname{tr}x=0\big\}.$
@@
for all $x\in\mathcal{O}$, see [17, (34.3)]. Together with Theorem 2.3(2), this implies that the subalgebra generated by an element $x$ coincides with the linear span of $x$ and xxx*x.
@@ 2.2 Multiplication table of Okubo algebras with isotropic norm
A special role in the study of Okubo algebras is played by algebras of the form $\mathcal{O}_{\alpha,\beta}$. Given nonzero $\alpha,\beta\in\mathbb{F}$, the multiplication table of $\mathcal{O}_{\alpha,\beta}$ is described by Table 1, while basis elements satisfy the conditions [15, pp. 3-4]:
@@
Theorem 2.7 ([11, Theorem 7], [15, Theorem 4.2]).
@@
Theorem 2.8 ([12, Theorem 3.18], [14, Theorem 5.9]).
@@
Definition 2.10 ([13, Definition 22], [14, Definition 6.2]).
@@
Theorem 2.11 ([4, Proposition 9.9 and Theorem 9.13], [13, Lemma 21], [14, Corollary 6.5]).
@@
The relationship between the quaternionic idempotent and all other idempotents (quadratic and singular) of the split Okubo algebra is described in [14, Proposition 8.7].
It is also known that if $\operatorname{char}\mathbb{F}=3$, then the norm on any Okubo algebra over $\mathbb{F}$ is isotropic [10, Theorem 5.1]. Besides, any Okubo algebra over a finite field is split [15, Theorem 3.2].
The following proposition utilizes the classification of Okubo algebras via central simple algebras of degree $3$ which was obtained in [7]. It should be noted that these results were originally stated for $\operatorname{char}\mathbb{F}\notin\{2,3\}$, however, they still remain valid for $\operatorname{char}\mathbb{F}=2$, see [10, pp. 286–287].
Proposition 2.12 ([7, p. 2486, p. 2501, Propositions 7.3 and 7.4], [9, Corollary 3.6]).
@@ 2.3 Zero divisors in Okubo algebras
  • The orthogonalizer of $a$ is $O_{\mathcal{A}}(a)=\big\{b\in\mathcal{A}\mid ab=ba=0\big\}$, i.e., the set of all elements in $\mathcal{A}$ which are orthogonal to $a$.
  • The left annihilator of $a$ is the set $l.\operatorname{Ann}_{\mathcal{A}}(a)=\big\{b\in\mathcal{A}\mid ba=0\big\}$.
  • The right annihilator of $a$ is the set $r.\operatorname{Ann}_{\mathcal{A}}(a)=\big\{b\in\mathcal{A}\mid ab=0\big\}$.
@@
$x^{\perp}=\{a\mid\operatorname{n}(x,a)=0\}.$
Lemma 2.17 ([18, Theorem 3.1, Proposition 3.4]).
@@
$l.\operatorname{Ann}_{\mathcal{O}}(x)$ $=x*\mathcal{O}=\{x*a\mid a\in\mathcal{O}\},$
$r.\operatorname{Ann}_{\mathcal{O}}(x)$ $=\mathcal{O}*x=\{a*x\mid a\in\mathcal{O}\}.$
Lemma 2.18 ([18, Proposition 3.8]).
@@
Note that [18, Proposition 3.7] also describes the intersection of two left (or two right) annihilators of linearly independent zero divisors in $\mathcal{O}$, but it will not be used in the current paper.
@@ 3 Orthogonality graph of an Okubo algebra
and thus $\operatorname{n}(x,y)=0$. Note that the desired equality can be obtained directly from [18, Proposition 3.7] by using the fact that $z\in(x*\mathcal{O})\cap(y*\mathcal{O})$. ∎
@@
$V^{\prime}_{x}=\big\{[y]=\mathbb{F}y\mid y\in O_{\mathcal{O}}(x)\big\}$
@@
$V=\big\{[x]=\mathbb{F}x\mid x\in Z(\mathcal{O}),\>\operatorname{n}(x,x*x)=0\big\}$
@@
$V=\big\{[x]=\mathbb{F}x\mid x\in Z(\mathcal{O}),\>\operatorname{n}(x,x*x)=0\big\}$
@@ 4 Number of shortest paths
$C_{(\mathcal{S},*,\operatorname{n})}(e)=\big\{x\in\mathcal{S}\mid e*x=x*e\big\},$
@@
$\operatorname{Fix}(\tau)=\big\{x\in\mathcal{S}\mid\tau(x)=x\big\}.$
Proposition 4.9 ([4, Proposition 9.9 and Theorem 9.13]).
@@
In [14, Theorem 6.3] all possible cases of Petersson algebras associated to the Zorn vector-matrix algebra over a field of characteristic $3$ are described in detail. We will only cite the first item of this statement.
Theorem 4.10 ([14, Theorem 6.3(i)]).
@@
Proposition 4.11 ([13, Proposition 24], [14, Theorem 8.5]).
@@
    $\big\{e+x\mid 0\neq x\in C_{\mathcal{O}}(e)\cap C_{\mathcal{O}}(e)^{\perp}\big\};$
@@
Let $\tau$ be the automorphism defined in Eq. (4.2) which corresponds to the idempotent $e$. Consider a new multiplication on $\mathcal{O}$ given by Eq. (4.3). Then, by Proposition 4.7, the Okubo algebra $(\mathcal{O},*,\operatorname{n})$ is the Petersson algebra associated to the Hurwitz algebra $(\mathcal{O},\cdot,\operatorname{n})$ with the unity $e$ and its automorphism $\tau$. It follows from [14, Theorem 6.3] that $\tau$ satisfies the conditions of Theorem 4.10. Then $(\mathcal{O},\cdot,\operatorname{n})$ is isomorphic to the Zorn vector-matrix algebra over $\mathbb{F}$, and there exists a canonical basis such that $\tau$ acts on basis vectors according to Eqs. (4.4). It remains to show that if xx=0x*x=0 for some $x\in\mathcal{O}$, then $x\in C_{\mathcal{O}}(e)=\operatorname{span}\{e_{1},e_{2},u_{1},u_{2},v_{1},v_{3}\}$.
@@
Consider $x=(z_{0,2}+z_{1,2}+z_{2,2})-(z_{0,1}+z_{2,1}+z_{1,1})$. Since $x\perp C_{\mathcal{O}}(e)$, the element $e+x$ is a singular idempotent (see also an example after Definition 22 in [13] and [12, Theorem 3.12]). It is shown in Lemma 3.14 that yy=0y*y=0 for all $y\in O_{\mathcal{O}}(x)$. Clearly, this property of the orthogonalizer is preserved under automorphism, so it holds for all elements of type (1).